Two friends are renting an apartment. They pay the landlord the first month's rent. The landlord also requires them to pay an additional half oa a month's rent for a security deposit. They total amount they pay the landlord before moving in is $1725. What is the monthly rent?
I think it would be x + 1/2x = 1725 but I'm not sure what x equals.

Answer by jim_thompson5910(35256)


x + 1/2x = 1725

2*(x+1/2x) = 2*1725

2x + x = 3450

3x = 3450

x = 3450/3

x = 1150

So the monthly rent is $1,150.